# EVA-MILP: Towards Standardized Evaluation of MILP Instance Generation

### v2

@@ Appendix I Solving Time Gap
The purpose of this experiment is to assess the similarity in computational hardness between two sets of MILP instances (a ’training’ set and a ’generated’ set) specifically concerning the time taken to solve them with the Gurobi optimizer under configured settings. The metric calculates the percentage difference between the average solving times of the two sets.
Methodology & Operations
  1. 1.
    Individual Instance Solving and Time Measurement:
    • Every MILP instance file within both the training and generated sets is individually subjected to the Gurobi optimization process.
    • Gurobi Solver Configuration:
      • Time Limit: A maximum solving duration of 300 seconds (solve.time_limit) is applied per instance. If an instance’s solve time reaches this limit, the recorded time will reflect this cap.
      • Threading: The Gurobi solver is configured to utilize 256 CPU threads (solve.threads) for processing each instance. The configuration uses a high thread count primarily to accelerate the overall experimental runtime. While absolute solve times are influenced by the number of threads, the focus of this metric is the relative difference in average solve times between the two instance sets. It is assumed that using the same high thread count consistently for both sets allows for a meaningful comparison of their relative hardness under these specific multi-threaded conditions.
    • Time Recording: The actual wall-clock time consumed by the core Gurobi optimization call for each instance is precisely measured and recorded. Instances that encounter errors during solving or do not yield a valid time measurement are excluded from subsequent calculations.
  2. 2.
    Relative Time Gap Calculation:
    • The "Solving Time Gap" is computed as the percentage difference between the two average times. The specific calculation used is:
@@
      Relative Time $\text{Relative Time Gap}=\frac{|\text{generated\_mean\_time}-\text{training\_mean\_time}|}{\max(\text{training\_mean\_time},10^{-10})}\times 100\%$
Key Parameters Influencing the Metric
  • Gurobi Time Limit per Instance: 300 seconds.
  • Gurobi Threads per Instance: 256.
This procedure results in a percentage value that indicates how similar the two sets of instances are in terms of their average solving time when using the specified Gurobi configuration (time limit and threads). A lower percentage signifies greater similarity in solver-conditional hardness.
I.1 Results
@@
We have omitted the G2MILP-generated Independent Set (IS) instances from this presentation. The rationale for this decision lies in the substantial surge in their computational complexity compared to the baseline instances, leading to a prohibitive increase in solution times by a factor of several thousands.
@@
Table 17: Solving Time Gap Comparison. The table shows the percentage difference in average solving time between instances generated by different models (with varying mask ratios $\eta$) and the original training set instances.
Training Set (Original Avg Time) Model $\eta$ Avg Time (s) Solving Time Gap (%)
Set Cover (0.2644s) G2MILP 0.10 0.3002 13.54
@@
Appendix J Hyperparameter Tuning on Gurobi
@@
The core goal is to automatically optimize a select set of Gurobi solver hyperparameters using the SMAC3 framework. The optimization aims to find a parameter configuration that minimizes the average wall-clock solving time for a specific collection of MILP instances designated as the "tuning set". Subsequently, the effectiveness of this optimized configuration is assessed by comparing its performance (average solve time) against Gurobi’s default settings on a separate, unseen "test set" of instances, thereby evaluating the generalization capability of the tuned parameters.
Methodology & Operations
  1. 1.
    Inputs & Configuration:
    • Tuning Instance Set: A collection of MILP instances used by SMAC3 to guide the optimization process.
    • Test Instance Set: An independent collection of MILP instances used for final performance evaluation.
    • Tuning Framework: SMAC3 (via the SMAC4HPO facade).
    • Target Solver: Gurobi.
  2. 2.
    Hyperparameter Space Definition:
    • The tuning process focuses on optimizing 8 key Gurobi parameters. The search space for each parameter is defined in gurobi_train.yaml under param_space:
      • Heuristics: Continuous float range [0.0, 1.0].
      • MIPFocus: Categorical integer choices {0, 1, 2, 3}.
      • VarBranch: Categorical integer choices {-1, 0, 1, 2, 3}.
      • BranchDir: Categorical integer choices {-1, 0, 1}.
      • Presolve: Categorical integer choices {-1, 0, 1, 2}.
      • PrePasses: Categorical integer choices {-1, 0, 1, …, 20}.
      • Cuts: Categorical integer choices {-1, 0, 1, 2, 3}.
      • Method: Categorical integer choices {-1, 0, 1, 2, 3, 4, 5}.
  3. 3.
    Automated Tuning via SMAC3:
    • Optimization Loop: SMAC executes an iterative Bayesian optimization process for a budget of 50 trials (tuning.n_trials). In each trial:
      • SMAC suggests a candidate Gurobi hyperparameter configuration (a specific set of values for the 8 parameters within their defined space).
      • Objective Function Evaluation: This suggested configuration is evaluated as follows:
          Every MILP instance within the tuning set (paths.custom_instances_dir) is solved using Gurobi configured with the candidate parameters.
          Each solve is subject to a strict time limit of 60 seconds (solve.time_limit) and utilizes 256 threads (solve.threads).
          The wall-clock time taken for each solve is recorded.
          The arithmetic mean of these solve times across all instances in the tuning set is calculated.
      • SMAC Feedback: This mean solve time is returned to SMAC as the performance "cost" of the evaluated configuration. SMAC uses this feedback (and results from previous trials) to update its internal models and decide which configuration to try next, aiming to minimize this mean solve time.
    • Result: After 50 trials, SMAC reports the configuration that achieved the lowest mean solve time on the tuning set. This is designated as the "best configuration".
  4. 4.
    Performance Evaluation on Test Set:
    • To assess how well the tuned parameters generalize, the performance of the "best configuration" found by SMAC is compared against Gurobi’s default settings. This comparison is performed on the independent test set (paths.test_instances_dir).
    • Procedure:
      • Each instance in the test set is solved once (evaluation.repeat: 1) using Gurobi with the default parameter settings (approximated for the 8 tuned parameters). The solve time is recorded, subject to the same 60s time limit and 256 thread count.
      • Each instance in the test set is then solved once using Gurobi with the "best configuration" found by SMAC, again subject to the 60s time limit and 256 threads. The solve time is recorded.
    • Calculation: The average solve time across all test set instances is calculated separately for the default configuration and the "best configuration".
  5. 5.
    Output & Metric:
    • The primary output is the comparison between the average solve time achieved by the default Gurobi settings and the SMAC-tuned "best configuration" on the test set.
    • This difference is often reported as a percentage improvement of the tuned configuration over the default.
    • The specific "best configuration" (the set of 8 hyperparameter values) found by SMAC is also a key output.
Key Parameters
  • SMAC Trials: 50 (tuning.n_trials).
  • Gurobi Time Limit (per solve): 60 seconds (solve.time_limit).
  • Gurobi Threads (per solve): 256 (solve.threads).
  • Evaluation Repetitions (per instance on test set): 1 (evaluation.repeat).
  • Tuning Instance Set Path: Value of paths.custom_instances_dir.
  • Test Instance Set Path: Value of paths.test_instances_dir.
@@
Table 18: Best Gurobi hyperparameter values found for each experiment.
@@
Column Abbreviations: Heur.: Heuristics, Focus: MIPFocus, VarBr.: VarBranch, BrDir.: BranchDir, Pres.: Presolve, PreP.: PrePasses, Cuts: Cuts, Meth.: Method.
Appendix K Initial Basis Prediction
@@
Following the approach of [36], this experiment aims to leverage Graph Neural Networks (GNNs) to predict a high-quality initial basis for the LP relaxation of MILP instances. The primary goal is to train a GNN model, typically on a specified dataset (e.g., generated instances), and evaluate whether using the initial basis predicted by this trained model can accelerate the MILP solving process in Gurobi compared to Gurobi’s default initialization. The evaluation measures performance improvements in terms of Gurobi runtime, node count, and iteration count on an unseen test set.
Methodology & Operations
The experiment follows a structured process encompassing data representation, feature engineering, GNN model training, basis generation/repair, and performance evaluation.
1. Problem Description and Data Representation
  • An MILP instance is represented by its constraint matrix $A\in\mathbb{R}^{m\times n}$, right-hand side $b\in\mathbb{R}^{m}$, objective coefficients $c\in\mathbb{R}^{n}$, and variable bounds $l^{x},u^{x}\in\mathbb{R}^{n}$.
  • The task focuses on predicting an initial basis for the LP relaxation, defined by a set of $m$ basic variables/slacks ($B_{x}\subset\{1..n\}$, $B_{s}\subset\{1..m\}$, $|B_{x}|+|B_{s}|=m$) such that the corresponding basis matrix is non-singular.
  • Each MILP instance is transformed into a bipartite graph $G=(V,E)$, where $V$ consists of $n$ variable nodes and $m$ constraint nodes, and an edge $(v_{i},w_{j})\in E$ exists if $A_{ji}\neq 0$, with the edge weight being $A_{ji}$.
2. Feature Engineering
  • Variable Node $v_{i}$ Features:
    1. 1.
      Objective coefficient $c_{i}$.
    2. 2.
      Variable density: $\text{nnz}(A_{:i})/m$.
    3. 3.
      Similarity to slack lower bounds: $\langle A_{:i},l^{s}\rangle/(\|A_{:i}\|\|l^{s}\|)$.
    4. 4.
      Similarity to slack upper bounds: $\langle A_{:i},u^{s}\rangle/(\|A_{:i}\|\|u^{s}\|)$.
    5. 5.
      Variable lower bound $l_{i}^{x}$ (0 if infinite).
    6. 6.
      Variable lower bound flag (-1 if $\infty$- ∞, 0 if finite).
    7. 7.
      Variable upper bound $u_{i}^{x}$ (0 if infinite).
    8. 8.
      Variable upper bound flag (1 if +$\infty$+ ∞, 0 if finite).
  • Constraint Node $w_{j}$ Features:
    1. 1.
      Similarity to objective coefficients: $\langle A_{j:},c\rangle/(\|A_{j:}\|\|c\|)$.
    2. 2.
      Constraint density: $\text{nnz}(A_{j:})/n$.
    3. 3.
      Similarity to variable lower bounds: $\langle A_{j:},l^{x}\rangle/(\|A_{j:}\|\|l^{x}\|)$.
    4. 4.
      Similarity to variable upper bounds: $\langle A_{j:},u^{x}\rangle/(\|A_{j:}\|\|u^{x}\|)$.
    5. 5.
      Constraint (slack) lower bound $l_{j}^{s}$ (0 if infinite).
    6. 6.
      Constraint lower bound flag (-1 if $\infty$- ∞, 0 if finite).
    7. 7.
      Constraint (slack) upper bound $u_{j}^{s}$ (0 if infinite).
    8. 8.
      Constraint upper bound flag (1 if +$\infty$+ ∞, 0 if finite).
  • Normalization: Continuous features (excluding flags) are standardized (z-score normalization) before being input to the model.
3. Graph Neural Network Model
  • Architecture: An InitialBasisGNN model with a bipartite graph structure is used. It consists of:
    • Initial MLP layers projecting input features (8 dims) to a hidden dimension (128, model.hidden_dim).
    • $L=3$ (model.num_layers) BipartiteMessagePassing layers that iteratively update variable and constraint node embeddings. Residual connections are used.
    • Separate output MLPs predicting 3-dimensional logits for variables and constraints.
    • Dropout (0.1, model.dropout) is applied for regularization.
  • Output and Knowledge Masking:
    • The output logits correspond to the three basis states: [NonbasicAtLower, Basic, NonbasicAtUpper].
    • Knowledge Masks are applied to the logits to ensure the model does not predict physically impossible basis statuses.
    • Softmax is applied to the masked logits to obtain probability distributions $p_{x_{i}},p_{s_{j}}$ over the three states for each variable and slack.
4. Model Training
  • Dataset: A single GNN model is trained using instances from a specified training directory (datasets.train_dir).
  • Training Loop:
    • The GNNTrainer orchestrates the training for a maximum of 800 epochs (training.epochs) using batches of size 32 (training.batch_size).
    • Loss Function: A WeightedCELoss is employed, incorporating:
        Label smoothing (0.1) to improve generalization.
        Class weights, computed per batch, to mitigate class imbalance.
    • Optimization: The Adam optimizer is used with an initial learning rate of 1e-3 (training.learning_rate) and L2 regularization (weight decay 1e-4, training.weight_decay).
    • Early Stopping: Training terminates early if the validation loss does not improve for 50 consecutive epochs (training.early_stopping).
  • Output: The model weights achieving the best performance on the validation set are saved to a file.
5. Initial Basis Generation and Repair
  • Candidate Selection: The $m$ variables/slacks with the highest predicted probability of being Basic are selected.
  • Basis Repair: An iterative process ensures the basis matrix is numerically stable by performing LU decomposition and replacing unstable columns.
  • Nonbasic State Encoding: Nonbasic variables/slacks are set to NonbasicAtLower or NonbasicAtUpper based on the GNN’s predictions.
6. Performance Evaluation
  • Setup: The evaluation uses an independent test set and the trained GNN model.
  • Procedure: Each test instance is solved twice using Gurobi:
    1. 1.
      Baseline: Gurobi’s default initialization.
    2. 2.
      Predicted: Gurobi is initialized using the repaired basis predicted by the GNN.
  • Metrics: Performance improvements in runtime, node count, and iteration count are calculated.
  • Output: Results are aggregated and saved to a JSON file, with optional visualizations.
Key Parameters
  • GNN Architecture: 3 layers, 128 hidden dim, 0.1 dropout.
  • Training: Max 800 epochs, Batch 32, LR 1e-3, Adam, Weighted CE Loss, Early Stopping 50, Seed 42.
  • Evaluation Gurobi: 600s Time Limit, 256 Threads, 0.01 MIP Gap.
  • Basis Repair: 1e-12 Pivot Threshold, 50 Max Iterations.
  • Datasets: Training (datasets.train_dir), Test (datasets.test_dir).
Table 19: GNN prediction initial basis: Runtime and performance improvement
Dataset Source/Model $\eta$ Default Time (s) Best Time (s) Improvement (%)
CA ACM-MILP 0.05 0.078 0.077 2.6%
0.10 0.079 0.076 2.6%
0.20 0.078 0.075 4.0%
DIG-MILP 0.05 0.079 0.076 3.5%
0.10 0.080 0.076 4.1%
Original Dataset 0.078 0.077 2.5%
IS ACM-MILP 0.10 0.236 0.239 -0.2%
G2MILP 0.05 0.237 0.240 -0.6%
Original Dataset 0.236 0.237 0.2%
SC G2MILP 0.01 0.236 0.221 10.6%
0.05 0.240 0.229 9.1%
0.10 0.242 0.224 12.4%
Original Dataset 0.236 0.226 9.1%
Appendix L Solver-internal features
The Experiment
The experiment consists of two main phases:
Phase 1: Data Generation - Solving Instances and Extracting Metrics
  1. 1.
    Objective: To solve each MILP instance in both the original and generated sets using Gurobi under identical, fixed settings, and to record detailed information about the solver’s internal operations from the execution logs.
  2. 2.
    Instance Processing (per instance in a set):
    • Fixed Solver Settings: The Gurobi solve is performed with consistent parameters specified in solver_info.yaml to ensure comparability:
      • Time Limit (solve.time_limit): 100 seconds.
      • Threads (solve.threads): 128.
      • MIP Gap Tolerance (solve.mip_gap): 1e-4.
      • Random Seed (solve.seed): 42 (for reproducibility).
      • Presolve (gurobi.presolve): -1 (Automatic).
      • MIPFocus (gurobi.mip_focus): 0 (Balanced).
      • Heuristics (gurobi.heuristics): 0.5 (Default).
      • Cuts (gurobi.cuts): -1 (Automatic).
      • Logging (gurobi.output_flag=1, gurobi.log_to_console=0): Enabled to generate detailed log files for each instance solve, but console output is suppressed. Additional logging parameters (log_file_append, diagnostics, display_interval) might be used for more detailed logs.
    • Heuristic Data Collection: The configuration (gurobi.use_callback: true) indicates that a Gurobi callback (HeuristicCallback, code in metrics.gurobi_callbacks) might be used during the solve. This callback aims to capture more detailed information about heuristic successes than might be available solely through standard log parsing.
  3. 3.
    Log Parsing & Metric Extraction:
    • After Gurobi finishes processing an instance (reaches optimality, time limit, etc.), its generated log file is parsed.
    • The following key metrics are extracted:
      • Root Node Gap: The relative duality gap calculated after solving the initial LP relaxation at the root node. Extracted by parsing specific lines in the log (e.g., Root relaxation: objective ... and Best objective ...).
      • Heuristic Success Count: Tallies the occurrences of various heuristic methods finding improved integer solutions. This uses regular expressions to find keywords (e.g., RINS, Feasibility Pump, Diving, FoundHeuristic, Improved solution) in the log. If the Gurobi callback was used successfully (heuristic_data_source="callback"), the counts collected by the callback might be prioritized; otherwise, counts derived from log parsing are used (heuristic_data_source="log_parsing"). Counts are stored per heuristic type (e.g., heur_RINS, heur_FoundHeuristic).
      • Cut Plane Usage: Parses the Cutting planes: section of the log file to count the number of times each specific type of cut (e.g., Gomory, Cover, MIR, FlowCover, ZeroHalf) was applied during the solve. These are stored as separate counts (e.g., cut_Gomory, cut_Cover).
    • Other potentially useful information like final status, solve time, node count, and iteration count are also extracted.
  4. 4.
    Output: For each dataset (original and generated), the extracted metrics for all its instances are aggregated and saved into a structured CSV file (e.g., solver_info.csv).
Phase 2: Comparative Analysis
  1. 1.
    Objective: To quantify the similarity between the original and generated instance sets based on the distributions of the solver metrics collected in Phase 1.
  2. 2.
    Input: The two CSV files generated in Phase 1 (one for original data, one for generated data). Paths are specified in the respective comparison config files (e.g., compare_rootgap.yaml).
  3. 3.
    Comparison Methodology (Metric-Specific):
    • Root Node Gap (compare_solver_rootgap.py):
      • The root_gap values are extracted from both CSV files.
      • The empirical distributions of these scalar values are compared using the 1-Wasserstein distance. A smaller distance indicates more similar distributions.
      • Statistical summaries and visualizations (e.g., histograms, KDE plots, CDFs, box plots) are generated.
    • Heuristic Success Count (compare_solver_heur.py):
      • The total count of success of heuristic (heur_FoundHeuristic) is selected for comparison.
      • The distributions of counts for this chosen heuristic are compared between the two datasets using the 1-Wasserstein distance.
      • Similar statistical summaries and visualizations are generated. (This process can be repeated for different heuristic types by changing the config. However, Gurobi may not provide detailed information about which specific heuristics were used. We recommend considering solvers that offer more detailed heuristic information for future experiments.)
    • Cut Plane Usage (compare_solver_cutplane.py):
      • The vectors of cut counts are extracted for each instance from both CSV files.
      • Preprocessing:
          Normalization: For each instance, the raw cut counts are converted into proportions (count of cut type / total cuts for that instance).
          Scaling: These proportions are then standardized (using StandardScaler by default, as per compare_cutplane.yaml) across both datasets combined to give features zero mean and unit variance.
      • Dimensionality Reduction (PCA): Principal Component Analysis is applied to the combined, scaled proportion data. The top 3 principal components (pca.n_components: 3) capturing the most variance are retained.
      • Comparison: The distributions of the instance scores along each of these top 3 principal components are compared between the two datasets using the 1-Wasserstein distance.
      • Statistical summaries, PCA variance explained, and visualizations (e.g., PCA scatter plots, score distributions, loading plots) are generated. The main focus is on the first principal component (PC1).
  4. 4.
    Output: The primary outputs are the Wasserstein distance values comparing the distributions for Root Node Gap, the selected Heuristic Success Count, and the dominant principal components of Cut Plane Usage. These distances serve as quantitative measures of similarity between the original and generated datasets in terms of these solver behaviors.
Key Parameters:
  • Gurobi Settings (Phase 1): Time Limit=100s, Threads=128, Seed=42, MIPGap=1e-4, plus specific Presolve/MIPFocus/Heuristics/Cuts settings.
  • Comparison Metric: 1-Wasserstein Distance.
  • Cut Plane Analysis: PCA with 3 components, Standard Scaling.
  • Heuristic Analysis: Comparison performed on a specific heuristic count column (e.g., heur_FoundHeuristic).
This detailed two-phase experiment first gathers data on how Gurobi interacts with each instance under fixed conditions, and then uses statistical techniques (e.g., Wasserstein distance, PCA) to compare the overall behavioral profiles of the original and generated instance sets.
@@ 6. Performance Evaluation
Principal_Component Wasserstein_Distance
@@
Table 24: Aggregated Cut Plane Counts (Part 1: Sum over 1000 instances)
Dataset Gomory ZeroHalf Clique MIR RLT FlowCover
IS (Raw) 1305 61590 89 58 43763 0
Combinatorial Auction (Raw) 11183 7014 8941 98 81 0
Set Cover (Raw) 1350 3473 80 5601 240 0
ACM-MILP CA ($\eta=0.1$) 6524 3132 26779 25 2 229
ACM-MILP IS ($\eta=0.1$) 1087 34032 33 13 22932 0
DIG-MILP CA ($\eta=0.05$) 6024 5629 462 531 114 0
G2MILP SetCover ($\eta=0.05$) 1330 3228 31 5454 209 0
Table 25: Aggregated Cut Plane Counts (Part 2: Sum over 1000 instances)
Dataset Cover ModK RelaxLift InfProof StrongCG ImplBound
IS (Raw) 0 0 0 0 0 0
Combinatorial Auction (Raw) 3 24 0 0 6 0
Set Cover (Raw) 0 8 0 0 0 1
ACM-MILP CA ($\eta=0.1$) 86 3 4 3 0 0
ACM-MILP IS ($\eta=0.1$) 0 1 0 0 0 0
DIG-MILP CA ($\eta=0.05$) 14 34 0 1 14 0
G2MILP SetCover ($\eta=0.05$) 0 8 0 0 0 0
L.1 Visualization

### v3

@@ Appendix H Solving Time Gap
This experiment assesses the similarity in computational hardness between a ’training’ set and a ’generated’ set of MILP instances. The metric is the relative percentage difference in their average solve times using the Gurobi optimizer.
Methodology: Each instance from both sets is solved under identical Gurobi configurations to ensure a fair comparison. The key parameters are:
  • Time Limit: 300 seconds per instance.
  • Threads: 256 per instance.
The average solve time is computed for each set. The final metric is calculated using the formula:
@@
$\text{Relative Time Gap}=\frac{|\text{generated\_mean\_time}-\text{training\_mean\_time}|}{\max(\text{training\_mean\_time},10^{-10})}\times 100\%$
A lower percentage signifies a greater similarity in computational hardness between the two instance sets.
H.1 Results
@@
The results are in Table 14. We have omitted the G2MILP-generated Independent Set (IS) instances from this presentation. The rationale for this decision lies in the substantial surge in their computational complexity compared to the baseline instances, leading to a prohibitive increase in solution times by a factor of several thousands.
@@
Table 14: Solving Time Gap Comparison. The table shows the percentage difference in average solving time between instances generated by different sources (with varying mask ratios $\eta$) and the original training set instances. Original refers to the baseline training set. G2MILP instances for IS reached the 300s time limit.
Dataset Source $\eta$ Avg Time (s) Solving Time Gap (%)
SC Original 0.2644
G2MILP 0.10 0.3002 13.54
@@
G2MILP 0.10 300.00 88815.23
0.05 300.00 88815.23
0.01 300.00 88815.23
Appendix I Hyperparameter Tuning for Gurobi
@@
The objective of this project is to automatically tune a key set of Gurobi solver hyperparameters using the SMAC3 framework. The goal is to find a parameter configuration that minimizes the average wall-clock solve time for a specific collection of MILP instances, designated as the tuning set. Subsequently, the performance of this optimized configuration is compared against Gurobi’s default settings on a separate, unseen test set to evaluate its generalization capability.
I.1 Methodology
The process is divided into five phases to ensure robust optimization and an unbiased evaluation.
Phase 1: Inputs and Setup
The process utilizes two independent sets of MILP instances: a tuning set, which is used to train the SMAC3 model and guide its search for optimal parameters, and a separate test set, which is held out exclusively for the final performance evaluation to ensure an objective assessment. The core tuning engine is SMAC3, which directly interfaces with the target solver, Gurobi.
Phase 2: Hyperparameter Space Definition
We selected eight critical Gurobi hyperparameters for optimization. The parameter space includes one continuous parameter, Heuristics, defined on the range [0.0, 1.0]. The remaining seven parameters are categorical: MIPFocus {0, 1, 2, 3}, VarBranch {-1, 0, 1, 2, 3}, BranchDir {-1, 0, 1}, Presolve {-1, 0, 1, 2}, PrePasses {-1, …, 20}, Cuts {-1, 0, 1, 2, 3}, and Method {-1, …, 5}.
Phase 3: Automated Tuning with SMAC3
SMAC3 performs an iterative optimization over a budget of 50 trials. Each trial begins with SMAC3 suggesting a new hyperparameter configuration from its internal model. Gurobi then uses this configuration to solve all instances in the tuning set, with each solve constrained to a 60-second time limit and using 256 threads. The resulting average solve time across the set is calculated and returned to SMAC3 as a cost metric. This feedback allows SMAC3 to update its model and inform its choice for the next trial. Upon completion, SMAC3 reports the best configuration found, which is the one that yielded the lowest average solve time.
Phase 4: Performance Evaluation on the Test Set
To assess generalization, we compare the best configuration found by SMAC3 against Gurobi’s default configuration on the independent test set. Both configurations are run under identical computational conditions (60-second time limit, 256 threads) to ensure a fair and direct comparison of their performance.
Phase 5: Outputs and Analysis
The final outputs synthesize the results into two key components. The primary metric is a comparison of the average solve times on the test set for the best versus default configurations, often expressed as a percentage improvement. The analysis also provides the best configuration itself, detailing the specific values for the eight optimized hyperparameters that yielded the top performance.
@@ I.1 Methodology
Table 15: Best Gurobi hyperparameter values found for each experiment. Column Abbreviations: Heur.: Heuristics, Focus: MIPFocus, VarBr.: VarBranch, BrDir.: BranchDir, Pres.: Presolve, PreP.: PrePasses, Cuts: Cuts, Meth.: Method.
@@
Appendix J Initial Basis Prediction
@@
Following (Fan and others, 2023), this experiment utilizes Graph Neural Networks (GNNs) to predict a high-quality initial basis for the LP relaxation of MILP instances. The objective is to train a GNN on **an augmented dataset (combining original and generated instances)** and evaluate whether its predicted basis accelerates Gurobi compared to default initialization. Performance is measured by Gurobi runtime, node count, and iteration count on an unseen test set.
J.1 Methodology & Operations
The methodology employs a six-phase approach: transforming MILP instances into graph representations, feature extraction, model training, and performance evaluation. This process ensures robust basis prediction while maintaining numerical stability.
Phase 1: Problem Description and Data Representation
MILP instances are defined by constraint matrix , RHS vector , objective , and variable bounds . The prediction task identifies an optimal initial basis, defined by sets and (where ), ensuring the basis matrix is non-singular.
Instances are transformed into bipartite graphs . contains variable and constraint nodes; represents structural relationships where edge exists if , weighted by .
Phase 2: Feature Engineering
Node representations capture both local structure and global characteristics. Variable node features (8 dimensions) include: objective coefficient , variable density , and similarity to slack bounds (e.g., ). Bounds are encoded via numerical values and binary flags (finite vs. infinite).
Constraint node features (8 dimensions) include: similarity to objective , constraint density , and similarities to variable bounds. Continuous features (excluding binary flags) undergo z-score standardization for training stability.
Phase 3: Graph Neural Network Model
The hierarchical GNN architecture processes the bipartite structure. An initial MLP projects 8-dimensional inputs to a 128-dimensional hidden space. The core comprises BipartiteMessagePassing layers with residual connections to aggregate neighborhood information.
Separate MLP heads output 3-dimensional logits for variables and constraints, corresponding to basis states: NonbasicAtLower, Basic, and NonbasicAtUpper. Knowledge masking enforces physical constraints by blocking impossible states based on bounds. Final probabilities are obtained via softmax on masked logits, utilizing dropout (0.1) for regularization.
Phase 4: Model Training
A GNNTrainer trains the model using **a composite dataset where generated instances are added to the original baseline**. The training runs for up to 800 epochs with a batch size of 32. The loss function uses label smoothing (0.1) to improve generalization and dynamic class weights to handle class imbalance. Optimization employs Adam (learning rate 1e-3) with L2 weight decay (1e-4). Early stopping terminates training if validation loss stagnates for 50 epochs, preserving the best-performing weights.
Phase 5: Initial Basis Generation and Repair
GNN predictions are converted into a numerically stable basis. Candidates are selected based on the highest predicted ”Basic” probabilities. To ensure non-singularity, a repair process performs LU decomposition on the candidate matrix, iteratively replacing unstable columns. Finally, nonbasic variables are assigned states (NonbasicAtLower or NonbasicAtUpper) according to their predicted probability distributions.
Phase 6: Performance Evaluation
The evaluation compares the GNN-predicted basis against Gurobi’s default initialization on an independent test set. Metrics include runtime, node count, and iteration count, providing a comprehensive assessment of the algorithmic efficiency gains achieved by the learned initialization.
Appendix K Solver-Internal Features
This experiment analyzes the behavioral patterns of the Gurobi solver to compare original and generated sets of MILP instances. The methodology involves a two-phase process: first, we gather detailed performance metrics by solving each instance, and second, we apply statistical techniques to quantify the similarity between the instance sets based on the solver’s behavior.
Methodology
Phase 1: Data Generation and Metric Extraction
In the first phase, we solve every MILP instance from both the original and generated collections using Gurobi under identical, standardized settings. This controlled environment ensures that any observed differences in performance are attributable to the instances themselves, not the solver’s configuration. We enable detailed logging to capture a comprehensive record of the solver’s internal operations.
From these logs, we systematically parse and extract key operational metrics. These metrics include the duality gap at the root node, which indicates initial problem difficulty; the success counts of various heuristics, which measure the effectiveness of the solver’s solution-finding strategies; and the usage frequency of different cut plane types (e.g., Gomory, Cover, MIR), which reveals the solver’s approach to tightening the LP relaxation. This process transforms the raw log files into structured data, ready for quantitative analysis.
Phase 2: Comparative Statistical Analysis
The second phase quantifies the similarity between the original and generated instance sets by comparing the distributions of the metrics collected in Phase 1. We primarily use the 1-Wasserstein distance as a robust measure of distributional difference—a smaller distance implies greater similarity.
For one-dimensional metrics like the root node gap and heuristic success counts, we directly compare their distributions. For the multi-dimensional cut plane data, we first normalize the usage counts for each instance into proportions. We then apply Principal Component Analysis (PCA) to reduce the dimensionality of this data, focusing the comparison on the most significant patterns of solver behavior. The Wasserstein distance is then used to compare the instance sets along these principal components. The final output consists of these distance values, which serve as a quantitative score of behavioral similarity between the two datasets.
@@
Principal_Component W-Dist.
@@
Table 20: Aggregated Cut Plane Counts (Sum over 1000 instances)
Dataset Gomory ZeroHalf Clique MIR RLT FlowCover Cover ModK RelaxLift InfProof StrongCG ImplBound
IS (Raw) 1305 61590 89 58 43763 0 0 0 0 0 0 0
Combinatorial Auction (Raw) 11183 7014 8941 98 81 0 3 24 0 0 6 0
Set Cover (Raw) 1350 3473 80 5601 240 0 0 8 0 0 0 1
ACM-MILP CA ($\eta=0.1$) 6524 3132 26779 25 2 229 86 3 4 3 0 0
ACM-MILP IS ($\eta=0.1$) 1087 34032 33 13 22932 0 0 1 0 0 0 0
DIG-MILP CA ($\eta=0.05$) 6024 5629 462 531 114 0 14 34 0 1 14 0
G2MILP SetCover ($\eta=0.05$) 1330 3228 31 5454 209 0 0 8 0 0 0 0
K.1 Visualization
Here are some figures 3 4,5,6 showing the results of solver internal features comparison.